# Climb: Class-imbalanced Learning Benchmark on Tabular Data

### v1

@@ 3 The Climb Benchmark
We compiled 73 naturally class-imbalanced tabular datasets provided by OpenML (Vanschoren et al., 2014) that span a wide real-world application domains with varying sizes and imbalance levels [^1] Access via: https://imbalanced-ensemble.readthedocs.io/en/latest/api/datasets. A statistical summary is provided in Figure 2. More detailed descriptions of each dataset can be found in Appendix A.1. They are selected using the following criteria:
@@ 3.3 Benchmarking Protocol
Given the strong performance and widespread use of tree-based models on tabular data and their close integration with certain CIL methods (e.g., Balanced Random Forest), we use decision trees as the base classifier to cooperate with all CIL algorithms. The ensemble size is set to 100 for all ensemble-based methods. To ensure fair and optimal evaluation, we perform hyperparameter tuning using Optuna (Akiba et al., 2019), with 100 optimization trials for each of the 23 CIL algorithms with tunable hyperparameters across all 73 datasets to determine the best-performing configurations. The search space and further details are provided in Appendix A.2.
@@ 4 Benchmark Results and Analysis
Following our rigorous benchmarking protocols, we conducted comprehensive experiments across all benchmark datasets to reveal insights into the classification performance, computational efficiency, and robustness of different CIL methods under varying levels of class imbalance. These experiments involved $\sim$0.8 million hyperparameter search trials, training of over 10 million base models, across 73 (datasets) $\times$ 30 (CIL methods) $\times$ 5 (splits) $=$ 10,950 dataset-method-split pairs.
@@ 4.1 Main Benchmark Results
We report the main benchmark results in Table 2. To better present insights from the large volume of numerical results, we grouped the 73 datasets by imbalance ratio (IR) into four categories: low (IR$<5$), medium (IR$\in[5,10)$), high (IR$\in[10,50)$), and extreme (IR$>50$) imbalance. We report the performance and ranking of each CIL method averaged over each dataset within each group.
Table 2: Main benchmark results. Given the large number of results, we group the 73 datasets by imbalance level into 4 categories and report the averaged AUPRC (AP), macro F1, and Balanced Accuracy (BAC) for each CIL method (in ×$\times 10^{-2}$). Detailed results for each dataset can be found in A.3. For a comprehensive evaluation, we also rank all methods on each dataset and metric, and report their average ranks. Color coding is used to show the performance gains (blue) or losses (red) relative to the base no-balancing method, with deeper colors indicating larger differences.
@@
Group Stat RUS CC IHT NM TL ENN RENN AKNN OSS NCR ROS SMT BSMT SSMT ASYN SPE BC BRF EE UBS UBA OBS SMBS OBA SMBA CS AdaC AdaBS AsyBS
IR $\in$ [0, 5) (28 datasets) Score ($\uparrow$) AP 51.0 49.4 48.0 45.8 45.5 51.5 53.6 53.4 52.9 51.7 53.7 51.1 51.6 51.8 52.1 51.6 59.3 57.8 57.7 59.0 58.6 59.0 52.9 53.7 58.4 58.9 51.2 52.5 52.4 52.5
@@
IR $\in$ [5, 10) (24 datasets) Score ($\uparrow$) AP 50.9 43.2 35.3 40.0 32.1 51.0 52.5 52.6 52.5 50.8 52.5 51.1 51.7 51.4 51.9 50.7 64.6 62.7 60.5 62.4 63.8 62.4 54.8 54.1 61.2 62.8 51.4 54.1 54.4 54.1
@@
IR $\in$ [10, 50) (15 datasets) Score ($\uparrow$) AP 34.9 23.6 17.4 27.1 14.5 35.1 36.2 36.2 35.9 35.5 36.3 34.1 34.3 35.8 35.0 34.1 47.1 41.4 38.4 41.9 44.9 41.9 36.7 36.6 45.0 46.0 34.1 36.4 36.7 36.4
@@
IR $\in$ [50, 1000) (6 datasets) Score ($\uparrow$) AP 42.6 18.9 15.9 33.0 13.5 44.2 45.0 44.1 44.9 44.3 44.6 41.7 37.1 42.2 41.9 34.3 57.5 50.1 32.9 35.5 43.3 35.5 45.0 40.5 56.4 56.0 41.7 48.1 46.6 48.1
@@ 4.2 Performance versus Runtime Analysis
Figure 3: Performance versus runtime analysis, following the dataset grouping in Table 2. The x-axis shows the average runtime of each CIL algorithm, and the y-axis shows the average AUPRC. Desired methods are closer to the upper-left corner with high accuracy and low computational cost. Different markers indicate different CIL method categories, the dashed line denotes the base model (no balancing) performance and runtime. More results with other metrics are in A.3.
@@ Conclusion.
Appendix A Appendix
This appendix provides additional details to support the experiments and findings presented in the main paper. Section A.1 presents details of the datasets used in our benchmark. Section A.2 describes further reproducibility details such as hyperparameter search strategy and runtime measurement protocols Finally, Section A.3 provides additional results, including further pairwise method win-ratio comparisons, and full per-dataset evaluation scores.
A.1 Datasets Details
@@
UCI_churn $3$ 18 5.9 Binary Business Customer churn prediction dataset with limited metadata.
@@
A.2 More Reproducibility Details
@@
[…]
@@
[…]
@@
A.3 Additional Results
Pairwise comparisons between all CIL methods.

### v2

@@ 3 The Climb Benchmark
We compiled 73 naturally class-imbalanced tabular datasets provided by OpenML (Vanschoren et al., 2014) that span a wide real-world application domains with varying sizes and imbalance levels [^1] Access via: https://imbalanced-ensemble.readthedocs.io/en/latest/api/datasets. A statistical summary is provided in Figure 2. More detailed descriptions of each dataset can be found in Appendix A. They are selected using the following criteria:
@@ 3.3 Benchmarking Protocol
Given the strong performance and widespread use of tree-based models on tabular data and their close integration with certain CIL methods (e.g., Balanced Random Forest), we use decision trees as the base classifier to cooperate with all CIL algorithms. The ensemble size is set to 100 for all ensemble-based methods. To ensure fair and optimal evaluation, we perform hyperparameter tuning using Optuna (Akiba et al., 2019), with 100 optimization trials for each of the 23 CIL algorithms with tunable hyperparameters across all 73 datasets to determine the best-performing configurations. The search space and further details are provided in Appendix B.
@@ 4 Benchmark Results and Analysis
Following our rigorous benchmarking protocols, we conducted comprehensive experiments across all benchmark datasets to reveal insights into the classification performance, computational efficiency, and robustness of different CIL methods under varying levels of class imbalance. These experiments involved $\sim$0.8 million hyperparameter search trials, training of over 10 million base models, across 73 (datasets) $\times$ 30 (CIL methods) $\times$ 5 (splits) == 10,950 dataset-method-split pairs.
@@ 4.1 Main Benchmark Results
We report the main benchmark results in Table 2. To better present insights from the large volume of numerical results, we grouped the 73 datasets by imbalance ratio (IR) into four categories: low (IR<5<5), medium (IR[$\in[5,10)$), high (IR[$\in[10,50)$), and extreme (IR>50>50) imbalance. We report the performance and ranking of each CIL method averaged over each dataset within each group.
Table 2: Main benchmark results. Given the large number of results, we group the 73 datasets by imbalance level into 4 categories and report the averaged AUPRC (AP), macro F1, and Balanced Accuracy (BAC) for each CIL method (in $\times 10^{-2}$). Detailed results for each dataset can be found in C. For a comprehensive evaluation, we also rank all methods on each dataset and metric, and report their average ranks. Color coding is used to show the performance gains (blue) or losses (red) relative to the base no-balancing method, with deeper colors indicating larger differences.
@@
Group Stat RUS $C$ IHT NM TL ENN RENN AKNN OSS NCR ROS SMT BSMT SSMT ASYN SPE BC BRF $E$ UBS UBA OBS SMBS OBA SMBA CS AdaC AdaBS AsyBS
IR $\in$ [0, 5) (28 datasets)
Score ($\uparrow$) AP 51.0 49.4 48.0 45.8 45.5 51.5 53.6 53.4 52.9 51.7 53.7 51.1 51.6 51.8 52.1 51.6 59.3 57.8 57.7 59.0 58.6 59.0 52.9 53.7 58.4 58.9 51.2 52.5 52.4 52.5
@@
IR $\in$ [5, 10) (24 datasets)
Score ($\uparrow$) AP 50.9 43.2 35.3 40.0 32.1 51.0 52.5 52.6 52.5 50.8 52.5 51.1 51.7 51.4 51.9 50.7 64.6 62.7 60.5 62.4 63.8 62.4 54.8 54.1 61.2 62.8 51.4 54.1 54.4 54.1
@@
IR $\in$ [10, 50) (15 datasets)
Score ($\uparrow$) AP 34.9 23.6 17.4 27.1 14.5 35.1 36.2 36.2 35.9 35.5 36.3 34.1 34.3 35.8 35.0 34.1 47.1 41.4 38.4 41.9 44.9 41.9 36.7 36.6 45.0 46.0 34.1 36.4 36.7 36.4
@@
IR $\in$ [50, 1000) (6 datasets)
Score ($\uparrow$) AP 42.6 18.9 15.9 33.0 13.5 44.2 45.0 44.1 44.9 44.3 44.6 41.7 37.1 42.2 41.9 34.3 57.5 50.1 32.9 35.5 43.3 35.5 45.0 40.5 56.4 56.0 41.7 48.1 46.6 48.1
@@ 4.2 Performance versus Runtime Analysis
Figure 3: Performance versus runtime analysis, following the dataset grouping in Table 2. The x-axis shows the average runtime of each CIL algorithm, and the y-axis shows the average AUPRC. Desired methods are closer to the upper-left corner with high accuracy and low computational cost. Different markers indicate different CIL method categories, the dashed line denotes the base model (no balancing) performance and runtime. More results with other metrics are in Appendix C.
@@ 4.3 Robustness Analysis
Additional results in the appendix.
Due to space limitations, we present the key results and insights in the main text. Appendix C includes results with hybrid sampling methods and GBDTs, pairwise win-ratio comparisons, full per-dataset evaluation scores, and runtime analyses.
@@ Conclusion.
Acknowledgments and Disclosure of Funding
This work is supported by the National Science Foundation (IIS-2117902 and 2433308), MIT-IBM Watson AI Lab, and IBM-Illinois Discovery Accelerator Institute. The views and conclusions are those of the authors and should not be interpreted as representing the official policies of the funding agencies or the government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes, notwithstanding any copyright notation here on.
Appendix
This appendix provides additional details to support the experiments and findings presented in the main paper. Section A presents details of the datasets used in our benchmark. Section B describes further reproducibility details such as hyperparameter search strategy and runtime measurement protocols Finally, Section C provides extended results and analyses, including additional CIL baselines (hybrid sampling methods and GBDTs), pairwise win-ratio comparisons, discussion on the advantages of Self-paced Ensemble, comparison with the BAF benchmark, and full per-dataset evaluation scores and runtime statistics.
Appendix A Datasets Details
@@
UCI_churn $33$ 18 5.9 Binary Business Customer churn prediction dataset with limited metadata.
@@
Appendix B More Reproducibility Details
@@
NearMiss $\texttt{n\_neighbors}\in[1,10]$
EditedNearestNeighbors $\texttt{n\_neighbors}\in[1,10]$, $\texttt{kind\_sel}\in\{\text{all},\text{mode}\}$
Repeated ENN $\texttt{n\_neighbors}\in[1,10]$, $\texttt{kind\_sel}\in\{\text{all},\text{mode}\}$
AllKNN $\texttt{n\_neighbors}\in[1,10]$, $\texttt{kind\_sel}\in\{\text{all},\text{mode}\}$
OneSideSelection $\texttt{n\_neighbors}\in[1,10]$
NeighborhoodCleaningRule $\texttt{n\_neighbors}\in[1,10]$, $\texttt{kind\_sel}\in\{\text{all},\text{mode}\}$, $\texttt{threshold\_cleaning}\in[0.0,1.0]$
SMOTE $\texttt{k\_neighbors}\in[1,10]$
BorderlineSMOTE $\texttt{k\_neighbors}\in[1,10]$, $\texttt{m\_neighbors}\in[1,10]$
SVMSMOTE $\texttt{k\_neighbors}\in[1,10]$, $\texttt{m\_neighbors}\in[1,10]$
ADASYN $\texttt{n\_neighbors}\in[1,10]$
SelfPacedEnsemble $\texttt{k\_bins}\in[1,10]$
@@
BalancedRandomForest $\texttt{max\_samples}\in[0.5,1.0]$, $\texttt{max\_features}\in[0.5,1.0]$
EasyEnsemble $\texttt{max\_samples}\in[0.5,1.0]$, $\texttt{max\_features}\in[0.5,1.0]$
RUSBoost $\texttt{learning\_rate}\in[0.0,1.0]$, $\texttt{algorithm}\in\{\text{SAMME},\text{SAMME.R}\}$
UnderBagging $\texttt{max\_samples}\in[0.5,1.0]$, $\texttt{max\_features}\in[0.5,1.0]$
OverBoost $\texttt{learning\_rate}\in[0.0,1.0]$, $\texttt{algorithm}\in\{\text{SAMME},\text{SAMME.R}\}$
OverBagging $\texttt{max\_samples}\in[0.5,1.0]$, $\texttt{max\_features}\in[0.5,1.0]$
SMOTEBoost $\texttt{learning\_rate}\in[0.0,1.0]$, $\texttt{algorithm}\in\{\text{SAMME},\text{SAMME.R}\}$, $\texttt{k\_neighbors}\in[1,10]$
SMOTEBagging $\texttt{max\_samples}\in[0.5,1.0]$, $\texttt{max\_features}\in[0.5,1.0]$, $\texttt{k\_neighbors}\in[1,10]$
AdaCost $\texttt{learning\_rate}\in[0.0,1.0]$, $\texttt{algorithm}\in\{\text{SAMME},\text{SAMME.R}\}$
AdaUBoost $\texttt{learning\_rate}\in[0.0,1.0]$, $\texttt{algorithm}\in\{\text{SAMME},\text{SAMME.R}\}$
AsymBoost $\texttt{learning\_rate}\in[0.0,1.0]$, $\texttt{algorithm}\in\{\text{SAMME},\text{SAMME.R}\}$
@@
Performance-runtime analysis with all metrics.
Appendix C Additional Experiments, Detailed Results, and Discussions
C.1 Results with Additional CIL methods
Table 5: Extended summary benchmark results with hybrid sampling methods (SMOTEENN, SMOTETomek) and GBDTs (XGBoost, LightGBM, CATBoost), this table extends the main results in Table 2 by including additional CIL baseline. Given the large number of results, we group the 73 datasets by imbalance level into 4 categories and report the averaged AUPRC (AP), macro F1, and Balanced Accuracy (BAC) for each CIL method (in $\times 10^{-2}$). Detailed results for each dataset can be found in C. For a comprehensive evaluation, we also rank all methods on each dataset and metric, and report their average ranks. Color coding is used to show the performance gains (blue) or losses (red) relative to the base no-balancing method, with deeper colors indicating larger differences.
Dataset Avg. Metric Base Undersample Cleaning Oversample Hybrid Undersample Ensemble Oversample Ensemble Cost-Sensitive GBDTs
Group Stat RUS $C$ IHT NM TL ENN RENN AKNN OSS NCR ROS SMT BSMT SSMT ASYN SENN STom SPE BC BRF $E$ UBS UBA OBS SMBS OBA SMBA AdaC AdaBS AsyBS CS XGB LGB CAT
IR $\in$ [0, 5) (28 datasets)
Score ($\uparrow$) AP 51.0 49.4 48.0 45.8 45.5 51.5 53.6 53.4 52.9 51.7 53.7 51.1 51.6 51.8 52.1 51.6 51.8 51.2 59.3 57.8 57.7 59.0 58.6 59.0 52.9 53.7 58.4 58.9 52.5 52.4 52.5 51.2 58.0 58.2 58.3
F1 72.0 70.4 67.1 66.0 65.5 72.8 73.9 73.3 73.0 73.0 74.2 72.4 72.9 72.9 73.2 72.8 72.6 72.5 77.9 76.8 76.7 78.0 77.5 78.0 73.7 74.3 76.5 77.4 73.2 73.2 73.2 72.3 76.4 76.3 75.7
BAC 72.1 73.3 71.7 73.8 69.7 73.3 76.5 76.7 76.4 73.6 76.6 72.5 73.5 73.6 73.8 73.6 75.5 73.1 78.5 77.3 79.7 79.7 78.7 79.7 73.7 74.8 75.7 77.0 73.2 73.1 73.2 72.4 75.8 75.9 75.5
Rank ($\downarrow$) AP 24.8 29.3 31.4 27.1 30.6 22.5 15.4 16.4 18.5 21.6 15.5 25.2 23.4 22.8 22.5 23.8 21.8 24.7 6.0 10.0 9.6 6.8 6.9 6.6 19.6 17.0 8.0 6.1 20.0 20.9 20.2 25.7 10.1 9.6 9.6
F1 24.1 30.4 32.0 33.0 32.0 21.2 16.4 19.1 21.1 19.9 16.0 24.2 22.5 21.8 20.6 22.3 22.2 23.3 5.0 8.9 10.2 5.7 6.7 5.6 19.0 15.7 10.2 6.6 19.1 20.8 19.3 24.9 9.9 9.4 10.9
BAC 28.3 24.0 28.1 22.7 29.5 22.8 13.2 12.7 14.2 21.4 13.1 27.5 23.5 22.5 20.9 22.4 17.1 24.2 7.1 9.9 3.9 3.9 6.2 3.7 23.0 17.2 14.4 10.1 22.9 25.0 23.5 28.7 14.0 13.5 14.9
IR $\in$ [5, 10) (24 datasets)
Score ($\uparrow$) AP 50.9 43.2 35.3 40.0 32.1 51.0 52.5 52.6 52.5 50.8 52.5 51.1 51.7 51.4 51.9 50.7 50.8 50.7 64.6 62.7 60.5 62.4 63.8 62.4 54.8 54.1 61.2 62.8 54.1 54.4 54.1 51.4 62.8 63.2 62.5
F1 74.7 68.4 56.7 64.2 57.4 74.6 75.1 75.0 75.0 74.5 75.2 74.4 75.1 75.0 75.3 74.6 73.6 74.6 79.7 78.7 76.7 78.7 79.3 78.7 75.6 75.8 77.0 78.5 75.4 75.4 75.4 74.5 78.2 78.3 77.3
BAC 74.7 76.0 71.2 76.5 70.8 74.9 77.6 78.0 77.7 74.8 77.8 74.4 76.4 75.8 76.4 75.7 77.8 76.0 82.1 80.8 83.1 82.9 82.5 82.9 75.4 76.8 75.6 77.7 75.0 75.1 75.0 74.5 77.0 77.0 76.0
Rank ($\downarrow$) AP 24.4 29.9 33.3 29.5 33.3 23.8 17.8 17.2 18.4 23.8 17.8 23.9 21.3 22.3 20.5 23.9 23.0 23.6 3.5 7.5 10.6 7.8 4.3 7.8 17.2 18.2 12.1 8.0 18.4 18.6 18.8 23.6 8.9 7.6 9.6
F1 22.5 32.0 34.2 33.3 34.2 21.9 18.7 18.6 18.6 22.2 18.3 22.2 19.7 21.0 18.3 22.4 24.7 21.9 4.5 7.9 15.4 8.1 4.5 7.8 16.9 16.8 13.3 7.2 17.4 18.1 17.8 21.9 8.7 8.6 10.5
BAC 26.3 21.6 29.3 20.0 29.5 25.2 16.4 15.4 16.1 24.5 16.1 26.5 17.7 21.5 17.6 22.0 14.9 19.4 5.9 7.5 3.2 3.9 3.7 3.2 22.1 16.7 20.6 13.5 24.0 24.2 24.3 26.1 16.5 16.4 18.2
IR $\in$ [10, 50) (15 datasets)
Score ($\uparrow$) AP 34.9 23.6 17.4 27.1 14.5 35.1 36.2 36.2 35.9 35.5 36.3 34.1 34.3 35.8 35.0 34.1 33.2 33.7 47.1 41.4 38.4 41.9 44.9 41.9 36.7 36.6 45.0 46.0 36.4 36.7 36.4 34.1 47.3 43.8 45.1
F1 61.6 51.2 35.8 52.9 35.6 61.6 62.3 61.9 61.5 61.9 62.4 61.0 61.0 62.1 61.7 61.0 59.5 60.3 65.5 62.1 59.7 62.9 64.0 62.9 61.6 61.7 64.0 65.8 61.5 61.6 61.5 61.0 65.7 63.0 63.6
BAC 61.8 63.6 56.7 65.8 54.3 62.1 64.5 65.0 64.3 62.3 64.1 60.8 63.5 63.7 63.5 63.2 65.2 63.0 70.7 67.8 72.2 72.1 70.4 72.1 61.1 64.0 62.4 65.1 61.5 61.2 61.5 61.0 64.1 61.6 62.1
Rank ($\downarrow$) AP 22.0 29.9 34.3 25.3 34.2 21.4 16.0 16.3 17.3 19.9 15.3 24.7 23.3 18.5 20.5 23.3 22.2 24.7 4.7 11.1 16.0 11.7 6.8 11.6 17.1 17.5 9.7 7.4 19.5 17.6 20.1 24.5 5.0 10.4 10.0
F1 17.0 32.8 34.4 31.7 34.2 17.7 12.7 14.5 16.9 15.7 13.7 21.6 21.4 14.5 17.6 21.3 25.3 23.3 5.7 16.7 22.7 14.0 9.0 13.7 17.2 17.9 12.1 6.5 18.1 17.5 18.1 21.3 6.7 13.7 13.1
BAC 23.1 19.2 31.6 13.7 32.3 22.5 16.7 15.1 17.5 21.7 16.7 27.7 17.0 15.3 17.1 19.9 13.9 18.7 5.6 7.8 2.9 2.8 5.7 2.3 26.4 16.3 20.7 15.5 24.1 26.3 24.9 26.7 15.9 23.4 22.6
IR $\in$ [50, 1000) (6 datasets)
Score ($\uparrow$) AP 42.6 18.9 15.9 33.0 13.5 44.2 45.0 44.1 44.9 44.3 44.6 41.7 37.1 42.2 41.9 34.3 34.7 34.4 57.5 50.1 32.9 35.5 43.3 35.5 45.0 40.5 56.4 56.0 48.1 46.6 48.1 41.7 58.9 48.9 58.1
F1 74.0 50.6 35.2 68.0 32.9 75.0 75.1 74.7 74.8 75.1 75.0 73.6 71.9 74.4 73.7 70.3 70.2 70.1 74.7 68.6 56.3 59.8 68.5 59.8 74.3 71.9 75.8 76.6 74.9 74.9 74.9 73.9 78.7 73.4 77.8
BAC 74.6 81.8 70.5 79.9 66.2 75.1 75.8 76.1 75.6 75.2 75.6 73.1 77.2 76.1 76.0 76.3 79.4 76.0 85.9 83.0 88.0 87.3 85.7 87.3 73.5 77.6 72.4 74.4 74.8 73.2 74.8 73.1 75.8 73.5 74.9
Rank ($\downarrow$) AP 19.2 32.7 34.2 24.0 34.8 19.5 14.8 16.0 15.7 16.7 17.8 22.0 21.0 17.8 19.5 25.2 26.2 24.3 4.5 10.8 21.0 18.8 12.7 18.5 16.7 19.3 9.8 9.0 13.0 16.2 13.3 21.3 6.7 10.2 6.8
F1 16.0 32.8 34.5 22.5 34.5 14.0 11.5 14.2 13.7 12.2 13.8 18.2 17.5 14.5 16.5 22.0 21.0 20.8 13.3 19.3 31.8 30.2 22.7 30.2 15.3 16.5 12.8 10.7 11.8 14.0 11.5 15.7 5.5 10.7 7.8
BAC 24.3 7.8 24.8 11.3 26.5 23.7 15.7 16.5 16.7 22.3 17.8 29.3 15.2 18.5 18.5 19.5 9.3 19.8 4.7 5.5 2.2 3.0 4.3 2.7 27.3 14.3 30.7 25.2 23.3 28.7 24.0 28.5 19.8 24.7 23.5
*Abbreviations: Random Undersampling (RUS), Cluster Centroids (CC), Instance Hardness Threshold (IHT), NearMiss (NM), Tomek Links (TL), Edited Nearest Neighbors (ENN), Repeated ENN (RENN), AllKNN (AKNN), One-Sided Selection (OSS), Neighborhood Cleaning Rule (NCR), Random Oversampling (ROS), SMOTE (SMT), Borderline SMOTE (BSMT), SVM SMOTE (SSMT), ADASYN (ASYN), SMOTEENN (SENN), SMOTE Tomek (STom), Self-paced Ensemble (SPE), Balance Cascade (BC), Balanced Random Forest (BRF), Easy Ensemble (EE), RUSBoost (UBS), UnderBagging (UBA), OverBoost (OBS), SMOTEBoost (SMBS), OverBagging (OBA), SMOTEBagging (SMBA), Cost-sensitive (CS), AdaCost (AdaC), AdaUBoost (AdaBS), AsymBoost (AsyBS), XGBoost (XGB), LightGBM (LGB), CATBoost (CAT).
Here, we provide additional benchmark results that incorporate hybrid sampling methods (SMOTEENN (Batista et al., 2004), SMOTETomek (Batista et al., 2003)) and popular gradient-boosted decision tree (GBDT) models (XGBoost (Chen and Guestrin, 2016), LightGBM (Ke et al., 2017), CatBoost (Hancock and Khoshgoftaar, 2020)). These methods were not discussed in detail in the main paper, as our primary focus is on establishing a fair and unified comparison of representative class-imbalanced learning (CIL) techniques under consistent experimental settings. Hybrid sampling methods tend to be slower, more complex, and often underperform relative to their simpler counterparts, while GBDT models rely on specialized base learners with optimization strategies that differ fundamentally from the scikit-learn trees used throughout our benchmark, making direct comparisons less meaningful. But still, we include these results here for completeness, as they were frequently raised during the review process and help further contextualize the scope and applicability of Climb.
For implementation, SMOTEENN and SMOTETomek are adopted from the imblearn (LemaÃŽtre et al., 2017) package with their default configurations, ensuring consistency with widely used practice. For GBDTs, we evaluate XGBoost, LightGBM, and CatBoost under the same ensemble size as the other ensemble-based CIL methods in the main paper. The extended results are summarized in Table 5. The latter full dataset-level results in Table 6-9 also include these new CIL methods.
Hybrid Sampling Methods.
The extended results highlight two consistent trends. First, the hybrid sampling methods SMOTEENN and SMOTETomek do not provide consistent benefits across imbalance levels. Their performance in terms of AP, F1, and BAC is typically comparable to or worse than their single-component counterparts (e.g., SMOTE or ENN/Tomek alone), and their average ranks remain relatively low. This confirms that the added complexity of combining oversampling and cleaning does not yield robust gains in practice.
Advanced GBDTs.
Second, the GBDT baselines (XGBoost, LightGBM, CatBoost) achieve strong overall results, often surpassing classical resampling-based methods, particularly under higher imbalance ratios. Nevertheless, they are not uniformly superior: ensemble-based CIL methods such as SPE, RUSBoost, and SMOTEBagging remain highly competitive, achieving comparable or better ranks in several imbalance groups. These findings indicate that while GBDTs constitute powerful baselines, well-designed CIL ensembles can match or exceed their performance, especially when tailored to severe imbalance scenarios.
C.2 Pairwise comparisons between all CIL methods.